# An accuracy-runtime trade-off comparison of scalable Gaussian process approximations for spatial data

### v1

@@ Abstract
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models widely employed in various fields such as spatial statistics, time series analysis, and machine learning. A drawback of Gaussian processes is their computational cost having $\mathcal{O}(N^{3})$ time and $\mathcal{O}(N^{2})$ memory complexity which makes them prohibitive for large datasets. Numerous approximation techniques have been proposed to address this limitation. In this work, we systematically compare the accuracy of different Gaussian process approximations concerning marginal likelihood evaluation, parameter estimation, and prediction taking into account the time required to achieve a certain accuracy. We analyze this trade-off between accuracy and runtime on multiple simulated and large-scale real-world datasets and find that Vecchia approximations consistently emerge as the most accurate in almost all experiments. However, for certain real-world data sets, low-rank inducing point-based methods, i.e., full-scale and modified predictive process approximations, can provide more accurate predictive distributions for extrapolation.
@@ 1 Introduction
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models which are widely used in spatial statistics (Cressie, 1993; Cressie and Wikle, 2011; Banerjee et al., 2014) and machine learning (Williams and Rasmussen, 2006). Estimation and prediction with Gaussian processes has $\mathcal{O}(N^{3})$ computational complexity and requires $\mathcal{O}(N^{2})$ memory. These costs render Gaussian processes computationally prohibitive, even when $N$ is only moderately large. To mitigate this problem, various approximations have been proposed; see Heaton et al. (2018) and Liu et al. (2019) for reviews. The goal of a Gaussian process approximation is to attain computational speed-up with as little degradation of quality as possible compared to exact calculations. Most approximations have tuning parameters which can be adjusted to balance a trade-off between accuracy and computational efficiency. In particular, many approximations recover the exact GP calculations in a limit case for a certain choice of tuning parameters. What matters in practice, though, is which approximation has the highest accuracy for a given computational budget. Analyzing the accuracy of Gaussian process approximations without taking computational cost into account thus provides only limited insights. The goal of this article is to compare various popular Gaussian process approximations regarding this accuracy-runtime trade-off. To the best our knowledge, no such comparison has been done before in a systematic manner. Heaton et al. (2018) also compare scalable methods for modeling spatial data. However, the runtimes of the different methods are vastly different and, what is more, the methods are run on very different computational environments. Similarly, Huang et al. (2021), Abdulah et al. (2022), and Hong et al. (2023) compare various methods for analyzing spatial data but do not consider runtimes in their analysis.
@@
We consider the following approximations: Vecchia approximations (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Katzfuss and Guinness, 2021), covariance tapering (Furrer et al., 2005), modified predictive process (MPP) approximations also called fully independent training conditional (FITC) approximations (Snelson and Ghahramani, 2005; Banerjee et al., 2008; Finley et al., 2009), full-scale approximations (FSA) combining predictive processes with tapering (Sang and Huang, 2012), multi-resolution approximations (MRA) (Katzfuss, 2017), the stochastic partial differential equations (SPDE)-based approximation introduced in Lindgren et al. (2011b), fixed rank kriging (FRK) (Cressie and Johannesson, 2008), and the Fourier transform-based periodic embedding approach of Guinness (2019). Note that we are mainly interested in understanding how well the different methods approximate a certain exact Gaussian process in a sense specified below. This means that we are not primarily interested in comparing different methods concerning pure prediction accuracy. If this were the goal, we would also have to include other methods such as tree-boosting, neural networks, and other machine learning models. Most of the methods we consider (specifically, Vecchia, MPP/FITC, FSA, SPDE, and MRA) do indeed aim to approximate an exact Gaussian process with a given covariance function. FRK and the periodic embedding approach, however, do not aim to approximate a certain Gaussian process. We nonetheless include these two methods in our comparison since they are frequently used in spatial statistics (Heaton et al., 2018).
@@
We compare the accuracy of the different approximations concerning log-likelihood evaluation, parameter estimation, and prediction on multiple simulated and real world datasets. For prediction accuracy, we distinguish between interpolation and extrapolation. For datasets with a moderate sample size, we also compare the quality of the approximations to exact calculations, i.e., when no approximation is used. For every approximation, we conduct multiple computations with varying complexity by adjusting tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots or inducing points for modified predictive process approximations. This allows us to assess the trade-off between the quality and the computational time of the different approximations. We find that Vecchia approximations consistently emerge as the most accurate in almost all experiments. However, for certain real-world data sets, low-rank inducing point-based methods, i.e., full-scale and modified predictive process approximations, can provide more accurate predictive distributions for extrapolation.
@@ 1.1 Background on Gaussian processes
A stochastic process $f(\boldsymbol{s}),s\in D\subset\mathbb{R}^{d},$ is a Gaussian process if $(f(\boldsymbol{s}_{1}),\dots,f(\boldsymbol{s}_{N}))$ follows a multivariate Gaussian distribution for any finite set of input locations $\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{N}\in D$. A Gaussian process is defined by a mean function $\mu(\cdot)$ and a positive definite covariance function $k(\cdot,\cdot)$:
$\begin{split}\mu(\boldsymbol{s})&=E[f(\boldsymbol{s})],\\ \quad k(\boldsymbol{s},\boldsymbol{s}^{\prime})&=E\left[(f(\boldsymbol{s})-\mu(\boldsymbol{s}))(f(\boldsymbol{s}^{\prime})-\mu(\boldsymbol{s}^{\prime}))\right],\quad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D.\end{split}$
In regression settings, one usually has a set of $N$ noisy observations $y_{1},\ldots,y_{N}$ at training locations $\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{N}$:
$y(\boldsymbol{s}_{i})=f(\boldsymbol{s}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2}_{n}),\quad i=1,\ldots,N.$
Assuming a zero prior mean and a parametric covariance function $k_{\theta}(\cdot,\cdot)$, where the covariance parameters denoted by $\theta$ also include the error term variance $\sigma^{2}_{n}$, the log-marginal likelihood is given by
@@ 2.1 Simulated data
$k_{M}\left(\boldsymbol{s},\boldsymbol{s}^{\prime}\right)=\sigma^{2}\left(1+\frac{\sqrt{3}\left\|\boldsymbol{s}-\boldsymbol{s}^{\prime}\right\|}{\rho}\right)\exp\left(-\frac{\sqrt{3}\left\|\boldsymbol{s}-\boldsymbol{s}^{\prime}\right\|}{\rho}\right),$
where $\sigma^{2}$ and $\rho$ are the marginal variance and range parameters to simulate data. We also add an independent Gaussian error term (“nugget effect”) with variance $\sigma_{n}^{2}$. The parameters are thus $\theta=\left(\sigma_{n}^{2},\sigma^{2},\rho\right)$. We simulate data using two different sample sizes $N\in\{10^{\prime}000;100^{\prime}000\}$ and three different range parameters $\rho\in\{\frac{0.05}{4.74};\frac{0.2}{4.74};\frac{0.5}{4.74}\}$. The division by $4.74$ is done for better interpretability of the range parameters as practical correlation ranges. The marginal and noise variances are set to $1$ and $0.5$ for all simulations. For every simulation scenario, we repeatedly simulate a certain number of datasets in order to analyze the variability of the evaluation metrics. The number of repetitions depends on the combination of the sample size and the specific task and is specified in Section 2.3.
@@ 2.2.1 House price data
The house dataset from the R package spData (Bivand et al., 2013) contains information about 25’357 single family houses sold in Lucas County, Ohio between 1993 and 1998. The response variable is the logarithmic selling price. We use the spatial coordinates as input locations for the Gaussian processes, and we assume that the mean is zero after centering the data using the sample mean of the training set. We do a random $70$/$30$ training-test split allocating $75117^{\prime}$ 751 observations to the training set, while the remaining data points are evenly divided between “interpolation” and “extrapolation” test sets each comprising $8033^{\prime}$ 803 data points. An illustration of the split as well as the distribution of house prices can be found in Figure 2.
@@ 2.2.2 Laegern satellite data
The Laegern dataset (Schneider et al., 2023) encompasses measurements of plant functional traits of the Laegern temperate mixed forest in Switzerland. We focus on the canopy height which is provided in scaled form between 00 and $1$. Values are recorded on a $1099\times 384$ grid with $6$ m spatial resolution. After removing missing values and excluding non-forest areas, there are $180279^{\prime}$ 180 observations left. Similarly to the housing dataset, we allocate 70% of the observations to the training set, 15% to the “interpolation” test set, and 15% to the “extrapolation” test set. A graphical illustration can be found in Figure 3.
@@ 2.2.3 MODIS 2016 satellite data
Next, we consider land surface temperatures (LST) recorded in degrees Celsius by the MODIS satellite (https://mrtweb.cr.usgs.gov/) on August 4, 2016. Measurements are made on a $500\times 300$ spatial grid with a resolution of $0.009273987$ degrees. The same data was used by Heaton et al. (2018), and we use the same train-test split as in Heaton et al. (2018) with $569105^{\prime}$ 569 observations in the training set and $74042^{\prime}$ 740 in the test set.Due to the presence of a clearly visible trend, see Figure 4, we include a linear regression term with longitude and latitude coordinates as well as an intercept term in the GP models.
@@ 2.2.4 MODIS 2023 satellite data
We additionally use another temperature data coming from the MODIS satellite recorded on August 20, 2023. In this case, the spatial grid has dimensions $1200\times 1200$ with a resolution of $926.6254$ meters. We use the same data split as in Gyger et al. (2024): out of the $000600^{\prime}$ 000 available observations, $000400^{\prime}$ 000 are used for the training set, while the remaining $000200^{\prime}$ 000 are allocated to the test set. Illustrations of the temperature data and the data split are shown in Figure 5. For the same reasons discussed previously, we again make use of the intercept and the coordinates as predictors.
@@ 2.3 Evaluation criteria
The following section introduces the criteria used to assess the quality of the different approximations. For each evaluation metric, we measure the wall-clock time needed to obtain the corresponding quantities using the computing environment presented in Section 2.5. For instance, when evaluating the accuracy of covariance parameter estimates, we record the time required to fit the models. In the simulation-based experiments, where multiple repetitions are conducted, both times and metrics are averaged over the performed iterations. We use a Matérn $3/2$ covariance function for all models where this applies, i.e., all GP approximations, except FRK and periodic embedding, as well as exact calculations. For the MODIS LST 2016 data set, we use an exponential covariance function since Heaton et al. (2018) also used this covariance function.
@@ 2.3.1 Log-likelihood evaluation
We first consider the accuracy of log-likelihood evaluations of the different GP approximation. In experiments with $N=10^{\prime}000$, we conduct $100$ repetitions and consider the absolute difference to the exact log-likelihood, i.e., the log-likelihood of the data-generating GP, as exact calculations can still be run in a reasonable amount of time. For the synthetic datasets with $N=100^{\prime}000$, we inspect the negative log-likelihood value itself by carrying out $20$ repetitions. For the simulated data, the log-likelihood of the training data is evaluated both at the true data-generating parameters and by using another set of parameters obtained by doubling each parameter value. For the real-world datasets, we evaluate the log-likelihood for all methods using the set of covariance parameters estimated by the most accurate method that could be implemented within a reasonable amount of time. This corresponds to using estimates of exact Gaussian processes for the housing data, and estimates of Vecchia approximations with the highest number of neighbors for the three other datasets. Additionally, we evaluate the log-likelihood at a second set of parameters obtained again by doubling the former parameters.
@@ 2.3.3 Prediction
We also evaluate the accuracy of approximate predictive distributions. On synthetic datasets, we use the true data-generating covariance parameters for prediction. For the real-world dataset, we use each technique’s estimated parameters. In the latter real-world experiments, prediction times include the time required to fit the model, as estimating the parameters is a necessary first step, whereas for the simulated dataset we only measure the time to calculate predictions. In the simulated setting, prediction is done for the latent process, and for the real-world datasets, we predict the response variable.
The first metric used is the RMSE between the predicted and the true values, given by:
$\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(f_{i}-\mu_{p,i}\right)^{2}}\quad\text{and}\quad\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(y_{i}-\mu_{p,i}\right)^{2}},$
@@
for the simulated and real-world data, respectively, where $N_{s}$ is the sample size of the considered test data, and $\mu_{p,i}$ indicates the predictive mean for the i-th observation. To asses the accuracy of probabilistic predictions, we consider the univariate log-score defined as average univariate Gaussian negative log-likelihood (NLL) of the latent and observable processes,
$\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(f_{i}-\mu_{p,i}\right)^{2}}{2\sigma_{p,i}^{2}}+\frac{1}{2}\log\left(2\pi\sigma_{p,i}^{2}\right)\right]\quad\text{and}\quad\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(y_{i}-\mu_{p,i}\right)^{2}}{2\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)}+\frac{1}{2}\log\left(2\pi\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)\right)\right],$
where $\sigma_{p,i}^{2}$ and $\hat{\sigma}_{n}^{2}$ indicate the predictive variance of the i-th latent process realization and the estimated noise variance. The former version for the latent process is used for the simulated data and the latter version for the real-world data.
@@
Lastly, in real-world applications, we also use the continuous ranked probability score (CRPS) (Gneiting and Raftery, 2007) which, in the case of Gaussian distributions, can be computed as:
$\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}\left(-\frac{1}{\sqrt{\pi}}+2\phi\left(\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)+\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\left(2\Phi\left(\frac{y_{p,i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)-1\right)\right),$
where $\phi$ and $\Phi$ denote the probability density function and the cumulative density function of a standard Gaussian variable.
@@
On datasets with moderate sample sizes, where exact Gaussian processes can be run in a reasonable amount of time, we additionally assess the quality of the approximations by comparing their predictions to the ones obtained using exact calculations. For this, we consider the RMSE of both approximate univariate predictive means and variances compared to the ones of exact GP calculations. Further, we consider the Kullback–Leibler (KL) divergence between the exact and approximate univariate predictive distributions. In the case of univariate Gaussian distributions, the KL divergence is given by
$KL(N\left(\mu_{q},\sigma_{q}\right),N\left(\mu_{p},\sigma_{p}\right))=\log\frac{\sigma_{p}}{\sigma_{q}}+\frac{\sigma_{q}^{2}+\left(\mu_{q}-\mu_{p}\right)^{2}}{2\sigma_{p}^{2}}-\frac{1}{2}.$
The number of simulation runs to evaluate the prediction accuracy for the simulated data is $50$ for $N=10^{\prime}000$ and $10$ for $N=100^{\prime}000$.
2.4 Gaussian process approximations and choice of tuning parameters

### v2

@@ Abstract
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models widely employed in various fields such as spatial statistics and machine learning. A drawback of Gaussian processes is their computational cost having $\mathcal{O}(N^{3})$ time and $\mathcal{O}(N^{2})$ memory complexity which makes them prohibitive for large data sets. Numerous approximation techniques have been proposed to address this limitation. In this work, we systematically compare the accuracy of different Gaussian process approximations concerning likelihood evaluation, parameter estimation, and prediction taking into account the computational time required to perform these tasks. In other words, we analyze the trade-off between accuracy and runtime on multiple simulated and large-scale real-world data sets. We find that Vecchia approximations consistently emerge as the most accurate in almost all experiments.
Keywords: Spatial statistics; Big data; Computational efficiency
@@ 1 Introduction
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models which are widely used in spatial statistics (Cressie and Wikle, 2011; Banerjee et al., 2014) and machine learning (Williams and Rasmussen, 2006). Estimation and prediction with Gaussian processes has $\mathcal{O}(N^{3})$ computational complexity and requires $\mathcal{O}(N^{2})$ memory. These costs render Gaussian processes computationally prohibitive, even when $N$ is only moderately large. To mitigate this problem, various approximations have been proposed; see Heaton et al. (2019) and Liu et al. (2020) for reviews. The goal of a Gaussian process approximation is to achieve computational efficiency with as little degradation of quality as possible compared to exact calculations. Most approximations have tuning parameters which can be adjusted to balance a trade-off between accuracy and computational efficiency. In particular, many approximations recover the exact GP calculations in a limit case for a certain choice of tuning parameters. What matters in practice, though, is which approximation has the highest accuracy for a given computational budget. Analyzing the accuracy of Gaussian process approximations without taking computational cost into account thus provides only limited insights. The goal of this article is to compare various popular Gaussian process approximations regarding this accuracy-runtime trade-off. To the best our knowledge, no such comparison has been done before in a systematic manner. Heaton et al. (2019) also compare scalable methods for modeling spatial data. However, the runtimes of the different methods are vastly different, and the methods are run using different numbers of CPU cores. Similarly, Huang et al. (2021), Abdulah et al. (2022), and Hong et al. (2023) compare various methods for analyzing spatial data but do not consider runtimes in their analysis.
@@
We consider the following approximations: Vecchia approximations (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Katzfuss and Guinness, 2021), covariance tapering (Furrer et al., 2005), modified predictive process approximations also called fully independent training conditional (FITC) approximations (Snelson and Ghahramani, 2005; Banerjee et al., 2008; Finley et al., 2009), full-scale approximations (FSA) combining predictive processes with tapering (Sang and Huang, 2012), multi-resolution approximations (MRA) (Katzfuss, 2017), the stochastic partial differential equations (SPDE)-based approximation introduced in Lindgren et al. (2011), fixed rank kriging (FRK) (Cressie and Johannesson, 2008), and the Fourier transform-based periodic embedding approach of Guinness (2019). Note that we are mainly interested in understanding how well the different methods approximate a certain exact Gaussian process in a sense specified below. This means that we are not primarily interested in comparing different methods concerning pure prediction accuracy. If this were the goal, we would also have to include other methods such as tree-boosting, neural networks, and other machine learning models. Most of the methods we consider (specifically, Vecchia, FITC, FSA, SPDE, and MRA) do indeed aim to approximate an exact Gaussian process with a given covariance function. FRK and the periodic embedding approach, however, do not aim to approximate a certain Gaussian process. We nonetheless include these two methods in our comparison since they are frequently used in spatial statistics (Heaton et al., 2019).
@@
We compare the accuracy of the different approximations concerning log-likelihood evaluation, parameter estimation, and prediction on several simulated and real world data sets. For prediction accuracy, we distinguish between interpolation and extrapolation. For data sets with a moderate sample size, we also compare the quality of the approximations to exact calculations, i.e., when no approximation is used. For every approximation, we conduct multiple computations with varying complexity by adjusting tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots or inducing points for modified predictive process approximations. This allows us to assess the trade-off between the quality and the computational time of the different approximations. We find that Vecchia approximations consistently emerge as the most accurate in almost all experiments.
@@ 1.1 Background on Gaussian processes
A stochastic process $f(\boldsymbol{s}),s\in D\subset\mathbb{R}^{d},$ is a Gaussian process if $(f(\boldsymbol{s}_{1}),\dots,f(\boldsymbol{s}_{N}))$ follows a multivariate Gaussian distribution for any finite set of input locations $\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{N}\in D$. A Gaussian process is defined by a mean function $\mu(\boldsymbol{s})=E[f(\boldsymbol{s})]$ and a positive definite covariance function $\quad k(\boldsymbol{s},\boldsymbol{s}^{\prime})=E\left[(f(\boldsymbol{s})-\mu(\boldsymbol{s}))(f(\boldsymbol{s}^{\prime})-\mu(\boldsymbol{s}^{\prime}))\right],\quad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$. In regression settings, one usually has a set of $N$ noisy observations $y_{1},\ldots,y_{N}$ at training locations $\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{N}$: $y(\boldsymbol{s}_{i})=f(\boldsymbol{s}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2}_{n}),\quad i=1,\ldots,N.$ Assuming a zero prior mean and a parametric covariance function $k_{\theta}(\cdot,\cdot)$, where the covariance parameters denoted by $\theta$ also include the error term variance $\sigma^{2}_{n}$, the log-marginal likelihood is given by
@@ 2.2.1 House price data
The house data set, available from the R package spData, contains information about 25’357 single family houses sold in Lucas County, Ohio between 1993 and 1998. The response variable is the logarithmic selling price. We use the spatial coordinates as input locations for the Gaussian processes, and we assume that the mean is zero after centering the data using the sample mean of the training set. We do a random $70$/$30$ training-test split allocating $75117^{\prime}$ 751 observations to the training set, while the remaining data points are evenly divided between “interpolation” and “extrapolation” test sets each comprising $8033^{\prime}$ 803 data points. An illustration of the split as well as the distribution of house prices can be found in Figure 12 in Appendix A.1.
@@ 2.2.2 Laegern satellite data
The Laegern data set (Schneider et al., 2023) encompasses measurements of plant functional traits of the Laegern temperate mixed forest in Switzerland. We focus on the canopy height which is provided in scaled form between 00 and $1$. Values are recorded on a $1099\times 384$ grid with $6$ m spatial resolution. After removing missing values and excluding non-forest areas, there are $180279^{\prime}$ 180 observations left. Similarly to the housing data set, we allocate 70% of the observations to the training set, 15% to the “interpolation” test set, and 15% to the “extrapolation” test set. A graphical illustration can be found in Figure 13 in Appendix A.1.
@@ 2.2.3 MODIS 2016 satellite data
Next, we consider land surface temperatures (LST) recorded in degrees Celsius by the MODIS satellite (https://mrtweb.cr.usgs.gov/) on August 4, 2016. Measurements are made on a $500\times 300$ spatial grid with a resolution of $0.009273987$ degrees. The same data was used by Heaton et al. (2019), and we use the same train-test split as in Heaton et al. (2019) with $569105^{\prime}$ 569 observations in the training set and $74042^{\prime}$ 740 in the test set.Due to the presence of a clearly visible trend, see Figure 14 in Appendix A.1, we include a linear regression term with longitude and latitude coordinates as well as an intercept term in the GP models.
@@ 2.2.4 MODIS 2023 satellite data
We additionally use another temperature data coming from the MODIS satellite recorded on August 20, 2023. In this case, the spatial grid has dimensions $1200\times 1200$ with a resolution of $926.6254$ meters. We use the same data split as in Gyger et al. (2024): out of the $000600^{\prime}$ 000 available observations, $000400^{\prime}$ 000 are used for the training set, while the remaining $000200^{\prime}$ 000 are allocated to the test set. Illustrations of the temperature data and the data split are shown in Figure 15 in Appendix A.1. For the same reasons discussed previously, we again make use of the intercept and the coordinates as predictors.
@@ 2.3 Evaluation criteria
The following section introduces the criteria used to assess the quality of the different approximations. For each evaluation metric, we measure the wall-clock time needed to obtain the corresponding quantities using the computing environment described in Section 2.5. For instance, when evaluating the accuracy of covariance parameter estimates, we record the time required to fit the models. In the simulation-based experiments, where multiple repetitions are conducted, both times and metrics are averaged over the performed iterations. We use a Matérn $3/2$ covariance function for all models where this applies, i.e., all GP approximations, except FRK and periodic embedding, as well as exact calculations. For the MODIS 2016 data set, we use an exponential covariance function since Heaton et al. (2019) also used this covariance function.
@@ 2.3.1 Log-likelihood evaluation
We first analyze the accuracy of log-likelihood evaluations of the different GP approximation. In experiments with $N=10^{\prime}000$, we perform $100$ simulation iterations and consider the absolute difference to the exact log-likelihood since exact calculations can still be run in a reasonable amount of time. For the synthetic data sets with $N=100^{\prime}000$, we inspect the negative log-likelihood value itself by carrying out $20$ repetitions. For the simulated data, the log-likelihood of the training data is evaluated both at the true data-generating parameters and by using another set of parameters obtained by doubling each parameter value. For the real-world data sets, we evaluate the log-likelihood for all methods using the set of covariance parameters estimated by the most accurate method that could be implemented within a reasonable amount of time. This corresponds to using estimates of exact Gaussian processes for the housing data, and estimates of Vecchia approximations with the highest number of neighbors for the three other data sets. Additionally, we evaluate the log-likelihood at a second set of parameters obtained again by doubling the former parameters.
@@ 2.3.3 Prediction
We also evaluate the accuracy of predictive distributions. On synthetic data sets, we use the true data-generating covariance parameters for prediction. For the real-world data set, we use each technique’s estimated parameters, and prediction times include the time required to fit the model, whereas for the simulated data set we only measure the time to calculate predictions. In the simulated setting, prediction is done for the latent process, and for the real-world data sets, we predict the response variable.
@@
The first metric used is the RMSE between the predicted and true values given by $\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(f_{i}-\mu_{p,i}\right)^{2}}$ and $\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(y_{i}-\mu_{p,i}\right)^{2}},$ for the simulated and real-world data, respectively, where $N_{s}$ is the sample size of the considered test data, and $\mu_{p,i}$ indicates the predictive mean for the i-th observation. To asses the accuracy of probabilistic predictions, we consider the univariate log-score defined as average univariate Gaussian negative log-likelihood (NLL) of the latent and observable processes, $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(f_{i}-\mu_{p,i}\right)^{2}}{2\sigma_{p,i}^{2}}+\frac{1}{2}\log\left(2\pi\sigma_{p,i}^{2}\right)\right]$ and $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(y_{i}-\mu_{p,i}\right)^{2}}{2\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)}+\frac{1}{2}\log\left(2\pi\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)\right)\right],$ where $\sigma_{p,i}^{2}$ and $\hat{\sigma}_{n}^{2}$ indicate the predictive variance of the i-th latent process realization and the estimated noise variance. The former version for the latent process is used for the simulated data and the latter version for the real-world data.
@@
In real-world applications, we also use the continuous ranked probability score (CRPS) (Gneiting and Raftery, 2007) which can be computed as $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}\left(-\frac{1}{\sqrt{\pi}}+2\phi\left(\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)+\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\left(2\Phi\left(\frac{y_{p,i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)-1\right)\right)$ Gaussian distributions, where $\phi$ and $\Phi$ denote the probability density function and the cumulative density function of a standard Gaussian distribution.
@@
On data sets with moderate sample sizes, we additionally assess the quality of the approximations by comparing their predictions to the ones obtained using exact calculations. For this, we consider the RMSE of both approximate univariate predictive means and variances compared to the ones of exact GP calculations. Further, we consider the Kullback–Leibler (KL) divergence between the exact and approximate univariate predictive distributions. In the case of univariate Gaussian distributions, the KL divergence is given by $KL(N\left(\mu_{q},\sigma_{q}\right),N\left(\mu_{p},\sigma_{p}\right))=\log\frac{\sigma_{p}}{\sigma_{q}}+\frac{\sigma_{q}^{2}+\left(\mu_{q}-\mu_{p}\right)^{2}}{2\sigma_{p}^{2}}-\frac{1}{2}.$ The number of simulation runs to evaluate the prediction accuracy for the simulated data is $50$ for $N=10^{\prime}000$ and $10$ for $N=100^{\prime}000$.
2.4 GP approximations and choice of tuning parameters